OpenStudy (anonymous):

the area of a rectangle is 65ft^2 and the length of the rectangle is double the width. find the dimensions

OpenStudy (anonymous):

Okay I'm going to try and explain this to you as best as I can. 65=lw that's your formula Area (65)= l(length) times w(width) it says your length is 2 times the width. let's name the width X if this is the case then the length =2x plug both those in the equation and you get 65= (x)(2x) which simplified equals 65=3x divide by 3 on both sides 65/3=3x/3 x= 21.6666... you can round to whatever it asks for Hope this helps!